# The First Search for Optical Transient as a Counterpart of a Month-timescale IceCube Neutrino Multiplet Event

### v1

@@ 1 Introduction
In this study, we perform archival search for an optical counterpart of an IceCube “triplet” event, ICT-MJD59015 identified by R. Abbasi et al. (2025). R. Abbasi et al. (2025) searches multiple neutrino detections with the time window of $T_{\rm w}=30~{}{\rm days}$ focusing on a relatively long time scale transients such as SNe and TDEs. For the optical counterpart search, we use ZTF archival data. By following a blind analysis approach, we determine the signal detection criteria without analyzing the data in the direction of the target IceCube triplet event. After establishing these criteria, we analyze the data in the direction to search for the optical counterpart. Based on the results, we perform a statistical analysis to give constraints to the properties of neutrino-emitting transients.
@@ 2.1 IceCube Data
As a whole dataset, the observed multiplets were consistent with background-only hypothesis, and the null observation provided a constraint on $\varepsilon_{\nu}$ and burst rates of transients. Nevertheless, the two most significant multiplets had false alarm rates (FARs) smaller than the inverse of the total live-time ($1/11.4~{}\mathrm{yr}^{-1}$), implying a possible excess. The directional uncertainty of the source candidates of the two multiplets were presented in the paper, and showed excellent localizations within a radius of $\sim 0.3^{\circ}$ at 90% containment. Of these two multiplets, ZTF data (explained later in Sec. 2.2) covered the localization area of the second most significant triplet (hereafter referred to as ICT-MJD59015) after the last neutrino detection. When a neutrino emission energy per transient source is $\varepsilon_{\nu}=2\times 10^{51}~{}\mathrm{erg}$ and their local volumetric event rate is $R_{0}=10^{-7}~{}\mathrm{Mpc}^{-3}\,\mathrm{yr}^{-1}$ (these choices are inferred with the cosmic neutrino background flux measured by IceCube, R. Abbasi et al. 2022), a signalness—a probability that ICT-MJD59015 was from astrophysical origin—is approximately 40%. Although it was pointed out that there was a 4FGL Fermi source in the uncertainty region of the triplet (S. Abdollahi et al., 2022; J. Ballet et al., 2023), the association was likely an accidental coincidence based on the inconsistency between IceCube sensitivity and the distance to the source ($z=0.46$). Table 1 summarizes the information of ICT-MJD59015  (R. Abbasi et al., 2025).
@@
Direction (RA, DEC) $=(0.58^{\circ}$, $-0.35^{\circ})$
@@ 2.2 ZTF data
By the ZTF data reduction pipeline (F. J. Masci et al., 2019), transient candidates are selected by performing image subtraction using the deep reference image. Then, sources with $>5\sigma$ detection in the subtracted image are publicly distributed through the alert stream (M. T. Patterson et al., 2019). Note that an “alert” is issued for each detection, and a certain “object” is associated with multiple “alerts”.
As the alerts in this stage are dominated by bogus detections mainly due to residual by imperfect image subtraction, pixels around bright stars, and cosmic ray hits, ZTF provides a recommended set of quality cut about the shape of the detected sources (see Section 3 for more details). Then, a further cut is applied to remove solar system (moving) objects by cross-matching with known asteroids and by imposing multiple, positive detections at the same position. Finally, to remove Galactic variable stars, the nature of the closest objects is evaluated based on the “star-galaxy separation score” (sgscore $=1$ means more star-like while sgscore $=0$ means more galaxy-like, Y. Tachibana & A. A. Miller 2018). An object detected in Pan-STARRS1 (PS1) image within 1.5 arcsec from the transient candidate is assigned as a host object. If the host object is a galaxy-like object with sgscore $<0.5$, such a candidate passes the selection.
@@ 3.2 Model of signal - optical emission from $\nu$ source
In this study, we choose $t_{\rm decay}$ and $L_{\rm pk}$ as primary parameters. For simplicity, we adopt $T=10^{4}$ K and $2\times 10^{4}$ K for superluminous SNe (SLSNe) and TDEs respectively, based on the typical blackbody temperatures of known SLSNe and TDEs (S. van Velzen et al., 2021; Y. Yao et al., 2023; Z. H. Chen et al., 2023). We examine two values of the peak time of the optical lightcurve: $\Delta t=0$ and 30 days, where $\Delta t=t_{\nu,1}-t_{\rm pk}$.
@@
[…]
@@
[…]
@@ 3.3 Background modeling
The empirical modeling of the background penetration is based on the control data samples constituted by the ZTF wide-field data. To this end, three samples are created for the purpose of investigating different selection processes. The Wide Field (WF) control data sample comprises observation data collected during a single epoch on July 30, 2020, in the $g$-band, encompassing a total of 126,645 ZTF alerts. The sky coverage is about $6500~{}\rm{deg^{2}}$ (see also Appendix A), excluding the localization area of ICT-MJD59015. The remaining two samples contain data collected during a series of observation epochs. These data are used to calculate the expected numbers of background and TPRs resulting from examining multiple observations. The Multiple Epoch control data sample (hereafter referred to as ME1) consists of data from 11 different epochs spanning from July 14 to July 31 of 2020, covering $300~{}\rm{deg^{2}}$. The ME2 sample is constituted by data from 6 epochs ranging from July 14 to July 23 with $750~{}\rm{deg^{2}}$ coverage. See Section 3.5 for more details about this multi-epoch analysis.
@@
[…]
$r$-band $0.578\pm{0.015}$- 0.123
[…]
$r$-band $0.995\pm{0.004}$- 0.123
level 3 selection (object-wise Galactic source cut) sgscore $0.995\pm{0.001}$ 1,809 $0.279\pm{0.007}$ $0.241\pm{0.017}$
Gaia variable $0.9987\pm{0.0006}$ 449 $0.069\pm{0.003}$ $0.111\pm{0.012}$
Gaia proper motion $0.9985\pm{0.0006}$ 416 $0.064\pm{0.003}$ $0.102\pm{0.011}$
AGN $0.9973\pm{0.0008}$ 397 $0.061\pm{0.003}$ $0.094\pm{0.011}$
level 4 selection (object-wise light curve cut) Duration ($\Delta t_{\rm neg}$) $0.9977\pm{0.0023}$ 138 $0.021\pm{0.002}$ $0.084\pm{0.010}$
Notes: For the level 1 and level 2 selections, we show the TPR for detections in typical magnitude range (18.5-19.0 mag). The uncertainties reported for the TPR give a 90% confidence interval. For the number of background objects, we show the cumulative number of alerts for the detection-wise criteria (level 1 and level 2) and cumulative number of objects for the object-wise criteria (level 3 and level 4) with the statistical uncertainties of ±$\pm{\sqrt{N}}$.
@@ 3.4 Event Selections
3.4.1 level 1 selection - detection-wise quality cut
@@
The level 1 selection is detection-wise quality cut to remove the bogus detections. For this selection, we impose the following criteria on the data quality and shape of detected sources as suggested by the ZTF team [^1] https://zwickytransientfacility.github.io/ztf-avro-alert/filtering.html: the number of bad pixel (nbad $=0$), the full width at half maximum (${\tt fwhm}<5~{}{\rm pixels}$), elongation (${\tt elong}\leq 1.2$), and the difference between aperture magnitude and point spread function (PSF)-fitted magnitude (abs(magdiff) $<0.1$). Additionally, the ZTF team developed a machine learning Real/Bogus (RB) classifier (A. Mahabal et al., 2019) which estimates the RB score for individual detection ranging from 0 (bogus-like) to 1 (real-like). We adopt the threshold for RB score suggested by ZTF (${\tt rb}\geq 0.65$).
@@
We estimate the TPR of this level 1 selection using the alerts of real transients. We use real transients reported by the Bright Transient Survey (BTS, C. Fremling et al., 2020; D. A. Perley et al., 2020), which is a magnitude-limited ($m<19$ mag in either the $g$- or $r$-band in peak) survey of extragalactic transients. For the BTS transient samples constructed through the BTS sample explorer [^2] https://sites.astro.caltech.edu/ztf/bts/explorer.php, we collect 22,733 alerts of the true samples within the period from May 2020 to August 2020. The TPR is estimated by the number of alerts passing the level 1 selection divided by the number of all the alerts associated with these BTS samples. Figure 3 shows TPR along with PSF magnitude of each alert. For a typical magnitude range ($18.5\leq\rm{PSFmag}\leq 19.0$), the TPR of the level 1 selection is $0.628\pm{0.017}/0.578\pm{0.015}$ in $g$/$r$-band with statistical uncertainties.
@@
We also estimate the number of background alerts penetrating the level 1 selection. Applying the selection to WF control data sample, 126,645 alerts are reduced to 33,339 alerts by the level 1 selection. This corresponds to a background rate of $5.13\pm 0.03$ alerts $\rm deg^{-2}$.
3.4.2 level 2 selection - detection-wise Galactic source cut
@@
The level 2 selection is detection-wise Galactic source cut to remove solar system objects and Galactic objects such as variable stars. The criteria applied in the standard ZTF alert flow, known as the “extragalactic cut,” are employed in this case as well: (1) no known solar system objects within 5 arcsec, (2) at least two $>5\sigma$ detections separated by $>30$ min, and (3) the detection with a positive flux in the subtracted image.
We estimate the TPR of the level 2 selection in the same way as the level 1 selection.
@@
However, the TPR against the criterion (2) is contingent upon the phase of the light curve. Consequently, we consider the contribution from this criterion at the subsequent stage when performing the statistical test on the neutrino transient source model (see Section 3.6). The TPR that successfully meets the level 2 selection criteria, excluding criterion (2), is determined using the same data-driven approach employed for the TPR calculation in the level 1 selection. We apply the level 2 cuts to the alerts of real BTS transients to probe the TPR.
@@
Figure 4 shows TPR as a function of PSF magnitude in each band. For a typical magnitude range ($18.5\leq\rm{PSFmag}\leq 19.0$), the TPR of level 2 selection is $0.948\pm{0.012}/0.995\pm{0.004}$ in $g$/$r$-band.
@@
3.4.3 level 3 selection - object-wise Galactic source cut
The level 3 selection is object-wise Galactic source cut for removal of remaining Galactic sources and AGNs. First, we basically follow the cut used in the ZTF filtering system: (1) if there is an object detected in Pan-STARRS1 (PS1) image within 1.5 arcsec, it should be a galaxy-like object based on “star-galaxy separation score” (sgscore: Y. Tachibana & A. A. Miller 2018) ranging from 0 (galaxy-like) to 1 (star-like). While ZTF applies a criterion of ${\tt sgscore}<0.5$, we instead adopt ${\tt sgscore}<0.83$, following the FoM (Figure of Merit) threshold described in Y. Tachibana & A. A. Miller (2018), in order to maintain a higher signal TPR. This threshold recover 99.5 % of galaxy-like objects, which corresponds to a TPR of $0.995$ in our filtering system.
This relaxed cut leads to an increase in the number of background objects passing this criterion. Therefore, we add further criteria using cross match with catalog sources as follows: (2) No Gaia variable sources within 1.5 arcsec (L. Rimoldini et al. 2023; L. Eyer et al. 2023, through Part 4 Variability catalog in Gaia Collaboration 2022a) [^3] Gaia variable sources catalog is constructed by the data taken well before the detection of ICT-MJD59015 (from July 25, 2014, to May 28, 2017). Thus, they are unlikely to be the source that emit neutrino flare with the short time duration ($T_{\rm w}=30~{}{\rm days}$) considered in this study., (3) No Gaia sources with a significant ($>3\sigma$) proper motion within 1.5 arcsec (Gaia Collaboration et al. 2023, through Part 1 Main source catalog in Gaia Collaboration 2022b), and (4) No known AGNs within 1.5 arcsec (M. P. Véron-Cetty & P. Véron, 2010; E. W. Flesch, 2015).
We estimate the fraction of our target lost by these cross matches. We use 9552 spectroscopically confirmed SNe from BTS. As a result, 12/14/26 SNe are removed by each criteria (Gaia variable/Gaia proper motion/AGN). From these, we estimate the TPR of each criterion as $0.9987\pm 0.0006$, $0.9985\pm 0.0006$, and $0.9973\pm 0.0008$ for the criteria (2), (3), and (4) above, respectively.
We note this TPR is subject to uncertainties especially if the signal is a TDE. TDEs occur at the center of host galaxies while majority of SNe in the BTS samples do not occur at the center of the galaxies. Thus, our criteria above may filter out a fraction of TDEs that occur AGN-like host galaxy, which result in a reduction of TPR. We have checked that this effect is not significant by using the known TDE samples, but the number of TDE samples is still small (71 samples in BTS), preventing the estimate with a high statistics.
@@
3.4.4 level 4 selection - object-wise light curve cut
Finally, we apply the level 4 selection, object-wise light curve cut to further remove remaining variable stars and AGNs. For this criterion, we use light curve of each object as additional information. For the light curve of each object obtained through the ZTF alert browker, ALeRCE API [^4] https://alerce.readthedocs.io/en/latest/index.html, we identify the significant ($>5\sigma$) negative detection. Negative detections mean that the fluxes significantly decrease compared with the fluxes in the reference images. We define $\Delta t_{\rm neg}$ as the time interval from the first to the last significant negative detection. This criterion of $\Delta t_{\rm neg}>50$ days effectively remove persistent objects such as variable stars and AGNs without removing SNe and TDEs.
For this criterion, we estimate the TPR using historical light curve data of BTS spectroscopically confirmed SNe. We use 891 SNe located within $z\leq 0.02$ from BTS samples with well-sampled light curves. Then, we remove 27 SNe with the reference images contaminated by SN fluxes. Among the remaining 864 SNe, only 2 SNe show the light curve with $\Delta t_{\rm neg}>50$ days. Thus, we estimate the TPR of the level 4 selection as $0.9977\pm{0.0023}$.

### v2

@@ 1 Introduction
In this study, we perform archival search for an optical counterpart of an IceCube “triplet” event (hereafter referred to as ICT-MJD59015) identified by R. Abbasi et al. (2025). R. Abbasi et al. (2025) searches multiple neutrino detections with the time window of $T_{\rm w}=30~{\rm days}$ focusing on a relatively long time scale transients such as SNe and TDEs. For the optical counterpart search, we use ZTF archival data. By following a blind analysis approach, we determine the signal detection criteria without analyzing the data in the direction of the target IceCube triplet event. After establishing these criteria, we analyze the data in the direction to search for the optical counterpart. Based on the results, we perform a statistical analysis to give constraints to the properties of neutrino-emitting transients.
@@ 2.1 IceCube Data
As a whole dataset, the observed multiplets were consistent with background-only hypothesis, and the null observation provided a constraint on $\varepsilon_{\nu}$ and burst rates of transients. Nevertheless, the two most significant multiplets had false alarm rates (FARs) smaller than the inverse of the total live-time ($1/11.4~\mathrm{yr}^{-1}$), implying a possible excess. The directional uncertainty of the source candidates of the two multiplets were presented in the paper, and showed excellent localizations within a radius of $\sim 0.3^{\circ}$ at 90% containment. Of these two multiplets, ZTF data (explained later in Sec. 2.2) covered the localization area of the second most significant triplet, ICT-MJD59015, after the last neutrino detection. From Figure 4 of R. Abbasi et al. 2025, when a neutrino emission energy per transient source is $\varepsilon_{\nu}=10^{52}~\mathrm{erg}$ and their local volumetric event rate is $R_{0}=10^{-8}~\mathrm{Mpc}^{-3}\,\mathrm{yr}^{-1}$ (these choices are inferred with the cosmic neutrino background flux measured by IceCube, R. Abbasi et al. 2022), a signalness—a probability that ICT-MJD59015 was from astrophysical origin—can be inferred to be approximately 60%. Here, signalness is defined as the fraction of the multiplet originating from the astrophysical transient source to the total number of multiplets with identical test statistics as the triplet in question. The test statistic is a metric that is employed to evaluate the astrophysical hypothesis in comparison to the atmospheric neutrino hypothesis. This evaluation is made under the assumption that the observed cosmic neutrinos are generated by transient sources (see Appendix A for the detail). Although it was pointed out that there was a 4FGL Fermi source in the uncertainty region of the triplet (S. Abdollahi et al., 2022; J. Ballet et al., 2023), the association was likely an accidental coincidence based on the inconsistency between IceCube sensitivity and the distance to the source ($z=0.46$). Table 1 summarizes the information of ICT-MJD59015  (R. Abbasi et al., 2025).
@@
Direction (RA, DEC) $(0.58^{\circ}$, $0.35^{\circ})$
@@ 2.2 ZTF data
By the ZTF data reduction pipeline (F. J. Masci et al., 2019), transient candidates are selected by performing image subtraction using the deep reference image. Then, sources with $5\sigma$ detection in the subtracted image are publicly distributed through the alert stream [^1] https://ztf.uw.edu/alerts/public/ (M. T. Patterson et al., 2019). Note that an “alert” is issued for each detection, and a certain “object” is associated with multiple “alerts”.
As the alerts in this stage are dominated by bogus detections mainly due to residual by imperfect image subtraction, pixels around bright stars, and cosmic ray hits, ZTF provides a recommended set of quality cut about the shape of the detected sources (see Section 3 for more details). Then, a further cut is applied to remove solar system (moving) objects by cross-matching with known asteroids and by imposing multiple, positive detections at the same position. Finally, to remove Galactic variable stars, the nature of the closest objects is evaluated based on the “star-galaxy separation score” (sgscore =1=1 means more star-like while sgscore =0=0 means more galaxy-like, Y. Tachibana & A. A. Miller 2018). An object detected in Pan-STARRS1 (PS1) image within 1.5 arcsec from the transient candidate is assigned as a host object. If the host object is a galaxy-like object with sgscore <0.5<0.5, such a candidate passes the selection.
@@ 3.2 Model of signal - optical emission from $\nu$ source
In this study, we choose $t_{\rm decay}$ and $L_{\rm pk}$ as primary parameters. For simplicity, we adopt $T=10^{4}$ K and $2\times 10^{4}$ K for superluminous SNe (SLSNe) and TDEs respectively, based on the typical blackbody temperatures of known SLSNe and TDEs (S. van Velzen et al., 2021; Y. Yao et al., 2023; Z. H. Chen et al., 2023). For SLSNe, we also tested other choice of temperatures and confirmed that the results remain largely unchanged as long as $T\gtrsim 5000~{\rm K}$. We examine two values of the peak time of the optical lightcurve: $\Delta t=0$ and 30 days, where $\Delta t=t_{\nu,1}-t_{\rm pk}$.
@@
$\varepsilon_{\nu}\sim\frac{12\pi H_{0}\int E_{\nu}\Phi_{\nu}dE_{\nu}}{c\xi_{z}R_{0}},$ (2)
where $\Phi_{\nu}$ is the cosmic muon neutrino background intensity (R. Abbasi et al., 2022), $\xi_{z}=H_{0}^{-1}\int dz(1+z)^{1-\gamma}(dt/dz)\psi(z)$ is the correction factor by the cosmological source density evolution, $\gamma$ is the neutrino spectral index, $\psi(z)$ is the redshift evolution of the source, $H_{0}$ is the Hubble constant, and $c$ is the speed of light. Given the redshift evolutions of TDEs (H. Sun et al., 2015; J. Necker et al., 2025) and SLSNe (e.g., S. Yoshida et al., 2022) with the observed neutrino spectral index $\gamma\simeq 2.3$, we obtain $\xi_{z}\sim 0.3$ and 2.5, respectively. We consider two cases for TDEs given the uncertainty of the event rate, and one case for SLSNe. The mean neutrino energy fluence is estimated to be $\varepsilon_{\nu}\simeq 1.5\times 10^{52}$ erg for TDE with $R_{0}=6\times 10^{-8}\rm~Mpc^{-3}~yr^{-1}$ (TDE-L; $R_{0}$ is given by the break in the luminosity function by Y. Yao et al. 2023), $\varepsilon_{\nu}\simeq 3\times 10^{51}$ erg for TDE with $R_{0}=3\times 10^{-7}\rm~Mpc^{-3}~yr^{-1}$ (TDE-F; the case where fainter TDE emit neutrinos), and $\varepsilon_{\nu}\sim 1.5\times 10^{51}$ erg for SLSNe with $R_{0}=1\times 10^{-7}\rm~Mpc^{-3}~yr^{-1}$ (SLSN; T. J. Moriya et al. 2019; C. Curtin et al. 2019).
@@
Using these values of $\varepsilon_{\nu}$ with the source redshift evolutions, we estimate the neutrino flux from each source and compute the Poisson probability of detecting it as a triplet event. By integrating over redshift, we calculate the number of triplet sources and the probability distribution functions of the sources of triplet events, $P_{\rm dist}(z)$ as shown in Figure 2 (see Eqs. (8) and (16) in S. Yoshida et al. 2022). Independently, we can evaluate the detection horizon of the optical emission from the neutrino source, $z_{\rm lim}$, given the values of $L_{\rm pk}$, $t_{\rm decay}$, $t_{\rm pk}$, $T$, and the sensitivity of ZTF. For luminous transients of $L_{\rm pk}\sim 10^{44}\rm~erg~s^{-1}$ that we are targeting in this study, $z_{\rm lim}\sim 0.4$ is achieved with ZTF. In this case, the probability that the source is located within $z_{\rm lim}$, i.e., $f(z<z_{\rm lim})=\int^{z_{\rm lim}}P_{\rm dist}(z)dz$, is estimated to be >0.999>0.999 for triplet events while it is $\sim 0.2$ for singlet events. In the case of non-detection of any astrophysical transient within $z_{\rm lim}$, this probability corresponds to the significance level to exclude the adopted parameters of the transient. The high $f(z<z_{\rm lim})$ demonstrates the advantage of follow-up of triplet events: thanks to the preference toward a lower redshift, a stronger constraint can be obtained as compared with singlet events.
@@ 3.3 Background modeling
The empirical modeling of the background penetration is based on the control data samples constituted by the ZTF wide-field data. To this end, three samples are created for the purpose of investigating different selection processes. The Wide Field (WF) control data sample comprises observation data collected during a single night on July 30, 2020, in the $g$-band, encompassing a total of 126,645 ZTF alerts. The sky coverage is about $6500~\rm{deg^{2}}$ (see also Appendix B), excluding the localization area of ICT-MJD59015. The remaining two samples contain data collected during a series of observation epochs. These data are used to calculate the expected numbers of background and TPRs resulting from examining multiple observations. The Multiple Epoch control data sample (hereafter referred to as ME1) consists of data from 11 different epochs spanning from July 14 to July 31 of 2020, covering $300~\rm{deg^{2}}$. The ME2 sample is constituted by data from 6 epochs ranging from July 14 to July 23 with $750~\rm{deg^{2}}$ coverage. We use all the area observed in these epochs except for Galactic plane (Galactic latitude $|b|<10~\rm{deg}$), whose characteristic are very different from those of the target area of the triplet event. See Section 3.5 for more details about this multi-epoch analysis.
@@
level 1 selection (removing bogus detections ) $g$-band $0.628\pm{0.010}$ 33,339 $5.13\pm{0.03}$3 - 0.123
$r$-band $0.578\pm{0.009}$ - 0.123
level 2 selection (removing moving objects ) $g$-band $0.948\pm{0.007}$ 11,622 $1.79\pm{0.02}$3 - 0.123
$r$-band $0.995\pm{0.003}$ - 0.123
level 3 selection (removing known variable sources) sgscore $0.9950\pm{0.0004}$ 1,809 $0.279\pm{0.007}$ $0.241\pm{0.017}$
Gaia variable $0.9987\pm{0.0004}$ 449 $0.069\pm{0.003}$ $0.111\pm{0.012}$
Gaia proper motion $0.9985\pm{0.0004}$ 416 $0.064\pm{0.003}$ $0.102\pm{0.011}$
AGN $0.9973\pm{0.0005}$} 397 $061\pm{0.003} 0.0$$094\pm{0.011}$
level 4 selection (removing variable sources using light curves) Duration ($\Delta t_{\rm neg}$) $0.9977\pm{0.0016}$} 138 $021\pm{0.002} 0.0$$084\pm{0.010}$
Notes: For the level 1 and level 2 selections, we show the TPR for detections in typical magnitude range (18.5-19.0 mag). The $1\sigma$ uncertainties reported for TPRs are calculated based on binomial distribution. For the number of background objects, we show the cumulative number of alerts for the detection-wise criteria (level 1 and level 2) and cumulative number of objects for the object-wise criteria (level 3 and level 4) with the $1\sigma$ statistical uncertainties ($\pm{\sqrt{N}}$ of the original numbers of background objects).
@@ 3.4 Event Selections
3.4.1 level 1 selection
@@
The level 1 selection is detection-wise cut to remove bogus detections. For this selection, we impose the following criteria on the data quality and shape of detected sources as suggested by the ZTF team [^2] https://zwickytransientfacility.github.io/ztf-avro-alert/filtering.html: the number of bad pixel (nbad =0=0), the full width at half maximum (${\tt fwhm}<5~{\rm pixels}$), elongation (${\tt elong}\leq 1.2$), and the difference between aperture magnitude and point spread function (PSF)-fitted magnitude (abs(magdiff) <0.1<0.1). Additionally, the ZTF team developed a machine learning Real/Bogus (RB) classifier (A. Mahabal et al., 2019) which estimates the RB score for individual detection ranging from 0 (bogus-like) to 1 (real-like). We adopt the threshold for RB score suggested by ZTF (${\tt rb}\geq 0.65$).
@@
We estimate the TPR of this level 1 selection using the alerts of real transients. We use real transients reported by the Bright Transient Survey (BTS, C. Fremling et al., 2020; D. A. Perley et al., 2020), which is a magnitude-limited ($m<19$ mag in either the $g$- or $r$-band in peak) survey of extragalactic transients. For the transients reported by BTS, we collected 22,733 associated alerts from May 2020 to August 2020. The alerts associated with BTS sample cover a wide magnitude range down to 21 mag. The TPR is estimated by the number of alerts passing the level 1 selection divided by the number of all the alerts associated with the transients reported by BTS. Figure 3 shows TPR along with PSF magnitude of each alert. For a typical magnitude range ($18.5\leq\rm{PSFmag}\leq 19.0$), the TPR of the level 1 selection is $0.628\pm{0.010}/0.578\pm{0.009}$ in gg/$r$-band ($1\sigma$ statistical uncertainties are calculated based on binomial distribution).
@@
We also estimate the number of background alerts penetrating the level 1 selection. Applying the selection to WF control data sample, 126,645 alerts are reduced to 33,339 alerts by the level 1 selection. This corresponds to a background rate of $5.13\pm 0.03$ alerts $\rm deg^{-2}$ (the error is based on the statistical uncertainties of $\pm{\sqrt{N}}$ for the original numbers of background objects, assuming Poisson statistics).
3.4.2 level 2 selection
@@
The level 2 selection is designed to remove moving objects in the Solar system while retaining candidates for new transients. The criteria applied in the standard ZTF alert flow are employed in this case as well: (1) no known solar system objects within 5 arcsec, (2) at least two $5\sigma$ detections separated by >30>30 min, and (3) the detection with a positive flux in the subtracted image.
@@
We estimate the TPR of the level 2 selection in the same way as the level 1 selection. However, the TPR against the criterion (2) is contingent upon the phase of the light curve. Consequently, we consider the contribution from this criterion at the subsequent stage when performing the statistical test on the neutrino transient source model (see Section 3.6) by imposing the criterion of multiple detections. The TPR that corresponds to the level 2 selection criteria, except criterion (2), is determined using the same data-driven approach employed for the TPR calculation in the level 1 selection. We apply the level 2 cuts to the alerts of real BTS transients to probe the TPR.
@@
Figure 4 shows TPR as a function of PSF magnitude in each band. For a typical magnitude range ($18.5\leq\rm{PSFmag}\leq 19.0$), the TPR of level 2 selection is $0.948\pm{0.007}/0.995\pm{0.003}$ in gg/$r$-band.
@@
3.4.3 level 3 selection
The level 3 selection is object-wise cut to remove remaining Galactic sources and AGNs. First, we basically follow the cut used in the ZTF filtering system: (1) if there is an object detected in Pan-STARRS1 (PS1) image within 1.5 arcsec, it should be a galaxy-like object based on “star-galaxy separation score” (sgscore: Y. Tachibana & A. A. Miller 2018) ranging from 0 (galaxy-like) to 1 (star-like). While ZTF applies a criterion of ${\tt sgscore}<0.5$, we instead adopt ${\tt sgscore}<0.83$, following the FoM (Figure of Merit) threshold described in Y. Tachibana & A. A. Miller (2018), in order to maintain a higher signal TPR. This threshold recover 99.5 % of galaxy-like objects, which corresponds to a TPR of $0.9950\pm{0.0004}$ in our filtering system.
This relaxed cut leads to an increase in the number of background objects passing this criterion. Therefore, we add further criteria using cross match with catalog sources as follows: (2) No Gaia variable sources within 1.5 arcsec (L. Rimoldini et al. 2023; L. Eyer et al. 2023, through Part 4 Variability catalog in Gaia Collaboration 2022a) [^3] Gaia variable sources catalog is constructed by the data taken well before the detection of ICT-MJD59015 (from July 25, 2014, to May 28, 2017). Thus, they are unlikely to be the source that emit neutrino flare with the short time duration ($T_{\rm w}=30~{\rm days}$) considered in this study., (3) No Gaia sources with a significant ($3\sigma$) proper motion within 1.5 arcsec (Gaia Collaboration et al. 2023, through Part 1 Main source catalog in Gaia Collaboration 2022b), and (4) No known AGNs within 1.5 arcsec (M. P. Véron-Cetty & P. Véron, 2010; E. W. Flesch, 2015).
We estimate the fraction of our target lost by these cross matches. We use 9552 spectroscopically confirmed SNe from BTS. As a result, 12/14/26 SNe are removed by each criteria (Gaia variable/Gaia proper motion/AGN). From these, we estimate the TPR of each criterion as $0.9987\pm{0.0004}$, $0.9985\pm{0.0004}$, and $0.9973\pm{0.0005}$ for the criteria (2), (3), and (4) above, respectively.
We note this TPR is subject to uncertainties especially if the signal is a TDE. TDEs occur at the center of host galaxies while majority of SNe in the BTS samples do not occur at the center of the galaxies. Thus, our criteria above may filter out a fraction of TDEs that occur AGN-like host galaxy, which result in a reduction of TPR. We have checked that this effect is not significant by using the known TDE samples, but the number of TDE samples is still small (71 samples in BTS), preventing the estimate with a high statistics. Also, two out of three TDEs reported to be associated with the IceCube neutrino detections occurred in AGN hosts. Our criteria above may exclude such TDEs occuring in AGNs.
@@
3.4.4 level 4 selection
Finally, we apply the level 4 selection, object-wise light curve cut ($\Delta t_{\rm neg}\leq 50~\rm{days}$) to further remove remaining variable stars and AGNs. For this criterion, we use light curve of each object as additional information. For the light curve of each object obtained through the ZTF alert broker, ALeRCE API [^4] https://alerce.readthedocs.io/en/latest/index.html, we identify the significant ($5\sigma$) negative detection. Negative detections mean that the fluxes significantly decrease compared with the fluxes in the reference images. We define $\Delta t_{\rm neg}$ as the time interval from the first to the last significant negative detection. This criterion of $\Delta t_{\rm neg}\leq 50$ days effectively remove persistent objects such as variable stars and AGNs without removing SNe and TDEs.
For this criterion, we estimate the TPR using historical light curve data of BTS spectroscopically confirmed SNe. We use 891 SNe located within $z\leq 0.02$ from BTS samples with well-sampled light curves. Before the TPR calculation, we remove 27 SNe with the reference images contaminated by SN fluxes. Among the remaining 864 SNe, only 2 SNe show the light curve with $\Delta t_{\rm neg}>50$ days. Thus, we estimate the TPR of the level 4 selection as $0.9977\pm{0.0016}$.
@@
Finally, we compare the TPR of level 3 and 4 selections with the corresponding TPR of the ZTF standard filtering system. In the ZTF standard filtering system, a stringent criterion of sgscore (${\tt sgscore}<0.5$ for PS1 objects within 1.5 arcsec) is imposed. To estimate the TPR of this criterion, we use transients reported in BTS between August 2019 and July 2020, which were selected using a looser condition of ${\tt sgscore}<0.76$ for PS1 objects within 2 arcsec (D. A. Perley et al., 2020). This estimate gives a TPR of 0.949. Note that this TPR should be regarded as an upper limit, since the true parent sample of transients would be larger.
On the other hand, in our filtering system, we impose a loose condition of sgscore (${\tt sgscore}<0.83$), whose TPR is robustly estimated to be 0.995 by Y. Tachibana & A. A. Miller (2018). An increase of background objects by this loose condition is further suppressed by additional criteria in level 3 and level 4 selections. The total TPR of our level 3 and 4 criteria is 0.987 ($0.995\times 0.9987\times 0.9985\times 0.9973\times 0.9977$, see Table 2). This demonstrates that our filtering system achieves a higher TPR than the ZTF standard filtering system.